OpenStudy (anonymous):

The radius of one sphere is twice as great as the radius of another sphere. Find the ratio of their surface areas.

Parth (parthkohli):

Why don't you take a radius of 1 and a radius of 2 and then compare their surface areas? :)

Parth (parthkohli):

\( \color{Black}{\Rightarrow Area(Sphere ) = 4\pi r^2 }\)

OpenStudy (anonymous):

I tried that. I got 1:4.

Parth (parthkohli):

\( \color{Black}{\Rightarrow \textbf{Radius of 1}: 4\pi r ^2 \Longrightarrow 4\pi }\) \( \color{Black}{\Rightarrow \textbf{Radius of 2}:4\pi r^2 \Longrightarrow 16\pi }\) 1:4 is correct~!

OpenStudy (anonymous):

Thanks!! :) :) :)